IF WE ARE GIVEN WITH TWO OR MORE DISPLACEMENT EQUATIONS OF ย  S.H.M OF FORM


y=Asin(wt+kx+phase angle)


THEN HOW TO FIND RESULTANT AMPLITUDE,PHASE ANGLE

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Aman

let

y1=A1sin(wt+kx)

y2=A2sin(wt+kx+Φ)

y=y1 +y2

  =A1sin(wt+kx)  +  A2sin(wt+kx+Φ)

  =A1sin(wt+kx )  +A2sin(wt+kx)cos(Φ) +A2cos(wt+kx)sin(Φ)

  =(A1 +A2 cosΦ  )sin(wt+kx ) +A2cos(wt+kx)sin(Φ)

   let  (A1 +A2 cosΦ  ) =A cosδ

and A2 sinΦ  =Asinδ

so y= A cosδ sin(wt+kx ) +Asinδcos(wt+kx)

        =Asin(wt+kx+δ )

  here A is resultant magnitude and  δ is phase angle

A=(A12 +A22 +2A1A2cosδ)1/2

tanδ =A2sinΦ/(A1 +A2cosΦ)

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

  

V.S.R Vijay
10 years ago

Dear Aman

let

y1=A1sin(wt+kx)

y2=A2sin(wt+kx+Φ)

y=y1 +y2

  =A1sin(wt+kx)  +  A2sin(wt+kx+Φ)

  =A1sin(wt+kx )  +A2sin(wt+kx)cos(Φ) +A2cos(wt+kx)sin(Φ)

  =(A1 +A2 cosΦ  )sin(wt+kx ) +A2cos(wt+kx)sin(Φ)

   let  (A1 +A2 cosΦ  ) =A cosδ i don''t understand why??

and A2 sinΦ  =Asinδ i don''t understand why?? please help me! 

so y= A cosδ  sin(wt+kx ) +Asinδcos(wt+kx)

        =Asin(wt+kx+δ )

  here A is resultant magnitude and  δ is phase angle

A=(A12 +A22 +2A1A2cosδ)1/2

tanδ =A2sinΦ/(A1 +A2cosΦ)